# STEMS: Spatial-Temporal Enhanced Safe Multi-Agent Coordination for Building Energy Management

### v1

@@ Abstract
Building energy management is essential for achieving carbon reduction goals, improving occupant comfort and reducing energy costs. Coordinated building energy management faces critical challenges in exploiting spatial-temporal dependencies while ensuring operational safety across multi-building systems. Current multi-building energy systems face three key challenges: insufficient spatial-temporal information exploitation, lack of rigorous safety guarantees, and system complexity. This paper proposes Spatial-Temporal Enhanced Safe Multi-Agent Coordination (STEMS), a novel safety-constrained multi-agent reinforcement learning framework for coordinated building energy management. STEMS integrates two core components: (1) a spatial-temporal graph representation learning framework using GCN-Transformer fusion architecture to capture inter-building relationships and temporal patterns, and (2) a safety-constrained multi-agent RL algorithm incorporating Control Barrier Functions to provide mathematical safety guarantees. Extensive experiments on real-world building datasets demonstrate STEMS’s superior performance over existing methods, showing that STEMS achieves 21% cost reduction, 18% emission reduction, and dramatically reduces safety violations from 35.1% to 5.6% while maintaining optimal comfort with only 0.13 discomfort proportion. The framework also demonstrates strong robustness during extreme weather conditions and maintains effectiveness across different building types.
@@ I Introduction
Buildings consume approximately 40% of global energy and account for 36% of CO2 emissions worldwide [1]. This massive energy consumption directly impacts environmental sustainability and economic costs as urbanization accelerates. Moreover, as urban residents spend over 80% of their time indoors, building energy management directly affects daily life quality. People require comfortable indoor environments under safe building control while facing rising energy costs and environmental concerns. This creates a complex multi-objective optimization challenge where comfort, safety, economic efficiency, and environmental sustainability must be balanced simultaneously [2]. Current building energy management systems operate independently among individual buildings, which limits their ability to achieve optimal performance [3].
@@
  • Insufficient spatial-temporal information exploitation: Current approaches fail to effectively exploit spatial-temporal information in building energy systems [4, 5]. As shown in Figure 1, buildings exhibit multiple types of relationships alongside their distinct energy consumption patterns. The spatial relationships include geographical proximity, attribute similarity, and temporal complementarity among different building types. Buildings close to each other will have heat transfer and shading effects among neighboring structures, residential buildings show double-peak patterns with morning and evening usage spikes, while commercial buildings maintain relatively stable consumption during business hours, However, existing methods ignore those hidden dependencies and patterns that are crucial for coordinated decision-making. Without proper spatial-temporal feature extraction and modeling mechanisms, buildings make decisions based on limited local observations, leading to suboptimal performance. For example, during summer peak hours, all buildings independently decide to pre-cool their spaces before peak pricing begins. This simultaneous cooling demand overloads the electrical grid, causing supply instability and higher costs for everyone.
@@
  • Lack of rigorous safety guarantees: Existing control methods lack rigorous safety guarantees [6, 7] or rely on soft constraints, such as penalty terms in RL algorithms, to discourage unsafe behaviors. However, these methods cannot provide mathematical guarantees against constraint violations [8]. During the learning and exploration process, control systems may still conduct unsafe actions such as exceeding battery capacity limits or unit power outage. Moreover, when buildings operate as isolated islands with limited local observations and broken information connections, they are more likely to exhibit safety warnings and uncoordinated consumption patterns, leading to power grid-level equipment damage risks and overloads, resulting in significant economic losses.
@@
  • System complexity and algorithmic challenges: The inherent complexity of building energy systems poses significant algorithmic challenges for coordination [9, 10]. As illustrated in Figure 1, the diverse building types and their complex interaction patterns create a challenging optimization landscape. Multi-objective optimization requirements involving comfort, efficiency, and safety lead to complex trade-offs that are difficult to balance simultaneously [11]. The rapid development of distributed storage and renewable energy systems creates dynamic and uncertain environments with intermittent generation and variable demand patterns. These factors, combined with the instability of RL algorithms in high-dimensional continuous spaces, make convergence slow and solution quality unpredictable [12, 13].
@@
To address these multi-agent RL challenges in building energy management, this paper proposes Spatial-Temporal Enhanced Safe Multi-Agent Coordination (STEMS), a novel spatial-temporal aware safe multi-agent reinforcement learning framework for coordinated building energy management. Our approach consists of two key components that work together to overcome the identified limitations. First, we develop a spatial-temporal graph representation learning framework that effectively captures the complex relationships among buildings. We propose a GCN-Transformer fusion architecture that processes both spatial dependencies through graph convolution and temporal patterns through attention mechanisms. This framework includes a selective information construction mechanism that enables intelligent collaboration among buildings by dynamically determining what information to share based on spatial-temporal contexts. The approach effectively handles high-dimensional state spaces through structured feature extraction and dimensionality reduction. Second, we design a safety-constrained multi-agent RL algorithm that provides mathematical guarantees for safe operation. We integrate Control Barrier Functions (CBFs) [8] as safety shields that ensure control actions remain within safe operating bounds. The algorithm incorporates a predictive safety checking mechanism that prevents unsafe actions from being executed while maintaining system performance. This enables precise safety control in continuous action spaces, addressing the critical need for reliable operation in real-world building energy systems. In summary, the main contributions of this work are threefold:
@@
  • We validate that STEMS significantly improves energy efficiency and economic performance through extensive experiments on real-world building energy datasets. The results demonstrate the effectiveness of our spatial-temporal aware approach and the importance of safety constraints in coordinated building energy management. STEMS achieves substantial improvements over existing approaches in both energy savings and cost reduction.
The remainder of this paper is organized as follows. Section II reviews related work in building energy management, spatial-temporal modeling, and safe multi-agent RL. Section III formulates the problem and presents our system model for coordinated building energy management. Section IV details our proposed spatial-temporal graph representation learning framework and safety-constrained multi-agent RL algorithm. Section V presents comprehensive experimental results and analysis on real-world building energy datasets. Section VI concludes the paper and discusses future research directions.
@@ II-A Problem Formulation
where $\mathcal{G}=(\mathcal{N},\mathcal{E})$ represents the inter-building communication network, $\mathcal{R}=\{R_{i}\}$ denotes the reward function set, $\mathcal{C}=\{C_{i}\}$ represents the constraint function set, and $\gamma$ is the discount factor.
@@ II-B Overall Framework
To simultaneously address spatial-temporal dependency modeling, safety constraint guarantees, and multi-objective optimization challenges in building energy management, STEMS adopts a modular design that integrates three synergistic core components: the spatial-temporal graph representation learning module, the CBF safety constraint checking module, and the multi-agent RL module. Figure 2 illustrates the overall architecture and information flow among these components. Initially, building agents collect comprehensive state observations $\mathcal{S}$ including internal device status, environmental conditions, and neighboring building information. These observations are then processed by the spatial-temporal graph representation learning module through observation embeddings, graph convolution and temporal attention layers to extract spatial-temporal features that capture inter-building relationships and temporal patterns. Subsequently multi-agent RL utilizes these enhanced spatial-temporal features to generate control actions $\mathcal{A}$ for each building through coordinated decision-making, where the rewards $\mathcal{R}$ and state transitions guide the continuous learning process.
@@
The renewable energy component maximizes clean energy utilization:
@@
These four components work together to achieve balanced performance. The economic term minimizes electricity costs by optimizing consumption timing and renewable energy use. For system stability, the grid-level coordination with smooth building operations helps buildings work together to avoid overloads while maintaining steady power consumption. The comfort term allows flexible trade-offs between energy efficiency and occupant comfort based on real-time conditions and the renewable component maximizes clean energy utilization across all buildings.
@@ II-D Safety-Constrained Decision Making
Building energy systems require strict safety constraints to prevent equipment damage and ensure reliable operation. Battery overcharging can cause thermal runaway and fire hazards. Power overloads can damage electrical equipment and cause grid instability.
Our safety-constrained approach ensures these critical requirements are never violated during both training and deployment. Unlike penalty-based methods that may still produce unsafe actions during exploration, we use CBFs to mathematically guarantee constraint satisfaction.
@@ II-D1 Control Barrier Function Framework
Building Safety: Building power consumption cannot exceed capacity to prevent equipment overload and local outages:
@@
The battery state will be updtaed with: $SOC_{t+1}^{i}=SOC_{t}^{i}+\frac{a_{battery}^{i}\cdot\Delta t}{\text{Capacity}_{i}}$.
@@
This ensures the executed action is as close as possible to the original policy output while maintaining safety. If it is feasible, the action is safe. Otherwise, emergency conservative actions will be applied.
@@
This approach provides mathematical guarantees that safety constraints are never violated, ensuring reliable building energy management operations.
@@ II-E Multi-Agent Training Algorithm
$\nabla_{\theta}J(\theta)=\mathbb{E}_{s,a\sim\mathcal{A}^{safe}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)\cdot A(s,a)\right]$ (24)
where $A(s,a)=R(s,a)+\gamma V_{\phi}(s^{\prime})-V_{\phi}(s)$ is the advantage function, and the expectation is taken only over the safe action space.
@@
$L(\phi)=\mathbb{E}_{s,a\sim\mathcal{A}^{safe}}\left[(R(s,a)+\gamma V_{\phi}(s^{\prime})-V_{\phi}(s))^{2}\right]$ (25)
@@
1:Initial policy parameters $\{\theta^{i}\}_{i=1}^{N}$, critic parameters $\{\phi^{i}\}_{i=1}^{N}$, graph parameters $\psi$
2:Trained policy networks $\{\pi_{\theta^{i}}^{*}\}_{i=1}^{N}$
@@
13:  Update actor networks using policy gradient: $\theta^{i}\leftarrow\theta^{i}+\alpha_{\pi}\nabla_{\theta^{i}}J(\theta^{i})$
14:  Update critic networks using TD error: $\phi^{i}\leftarrow\phi^{i}+\alpha_{v}\nabla_{\phi^{i}}L(\phi^{i})$
15:  Update graph parameters end-to-end: $\psi\leftarrow\psi+\alpha_{g}\nabla_{\psi}L_{total}$
16:  if episode $e$ mod $K=0$ then
17:   Adjust safety constraint parameters based on system performance
18:  end if
19:end for
20:return Trained policy networks $\{\pi_{\theta^{i}}^{*}\}_{i=1}^{N}$
@@ III-A1 Simulation Environment
We conduct experiments based on the CityLearn environment[14], a multi-building energy management simulation platform that provides realistic building energy dynamics and grid interactions. The experimental dataset is derived from real building energy consumption data in Travis County, Texas. The dataset spans from August 2018 to August 2019, capturing a full year of operation under typical subtropical humid climate conditions: annual average temperature of 20.7°C, temperature range from -7.8°C to 42.8°C. This climate profile creates significant cooling-dominated energy demands during summer months and moderate heating requirements in winter, providing realistic conditions for evaluating multi-building energy coordination strategies. The data includes 1 hour interval measurements of electricity consumption, solar generation, battery storage, and environmental conditions. The experimental setup consists of 8 buildings in a neighborhood configuration with typical regional distribution (5 residential, 2 commercial, 1 mixed-use). The 8 buildings are randomly selected from the Travis County dataset based on above ratio to ensure representative diversity in building characteristics and energy consumption patterns. Most residential buildings show dual-peak patterns with morning (6-9h) and evening (18-22h) usage spikes, reflecting typical household activities with strong seasonal variations. Commercial buildings demonstrate business-hour concentrated loads (9-17h) with midday peaks around 12-13h. Mixed-use buildings combine both residential and commercial characteristics with variable all-day consumption patterns. Each building is equipped with controllable devices including battery energy storage systems (BESS), HVAC systems, solar photovoltaic panels, hot water tank. Buildings can communicate with each other to share state information and coordinate energy management decisions.
@@ III-A2 Baseline Methods
Single-Agent SAC: Independent Soft Actor-Critic reinforcement learning for each building without coordination[15].
MADDPG: Multi-Agent Deep Deterministic Policy Gradient with centralized training and decentralized execution[16].
MetaEMS: Meta-learning based energy management system using Model-Agnostic Meta-Learning (MAML) framework[17].
MARLISA: Multi-Agent Reinforcement Learning with Independent Soft Actor-Critic operating independently[12].

### v2

@@ Abstract
Building energy management is essential for achieving carbon reduction goals, improving occupant comfort and reducing energy costs. Coordinated building energy management faces critical challenges in exploiting spatial-temporal dependencies while ensuring operational safety across multi-building systems. Current multi-building energy systems face three key challenges: insufficient spatial-temporal information exploitation, lack of rigorous safety guarantees, and system complexity. This paper proposes Spatial-Temporal Enhanced Safe Multi-Agent Coordination (STEMS), a novel safety-constrained multi-agent reinforcement learning framework for coordinated building energy management. STEMS integrates two core components: (1) a spatial-temporal graph representation learning framework using GCN-Transformer fusion architecture to capture inter-building relationships and temporal patterns, and (2) a safety-constrained multi-agent RL algorithm incorporating Control Barrier Functions to provide mathematical safety guarantees. Extensive experiments on real-world building datasets demonstrate STEMS’s strong performance over existing methods, showing that STEMS achieves 21% cost reduction and 18% emission reduction over traditional methods, 65% safety violation reduction compared with learning-based baselines, and maintains the lowest discomfort rate. The framework also demonstrates strong robustness during extreme weather conditions and maintains effectiveness across different building types.
@@ I Introduction
Buildings consume approximately 40% of global energy and account for 36% of CO2 emissions worldwide [chen2024green]. This massive energy consumption directly impacts environmental sustainability and economic costs as urbanization accelerates. Moreover, as urban residents spend over 80% of their time indoors, building energy management directly affects daily life quality. People require comfortable indoor environments under safe building control while facing rising energy costs and environmental concerns. This creates a complex multi-objective optimization challenge where comfort, safety, economic efficiency, and environmental sustainability must be balanced simultaneously [zhang2022building-survey]. Current building energy management systems operate independently among individual buildings, which limits their ability to achieve optimal performance [kumari2024multi].
@@
  • Insufficient spatial-temporal information exploitation: Current approaches fail to effectively exploit spatial-temporal information in building energy systems [haidar2023selective, chen2023graph]. As shown in Figure 1, buildings exhibit multiple types of relationships alongside their distinct energy consumption patterns. The spatial relationships include geographical proximity, attribute similarity, and temporal complementarity among different building types. Buildings close to each other will have heat transfer and shading effects among neighboring structures, residential buildings show double-peak patterns with morning and evening usage spikes, while commercial buildings maintain relatively stable consumption during business hours. However, existing methods ignore these hidden dependencies and patterns that are crucial for coordinated decision-making. Without proper spatial-temporal feature extraction and modeling mechanisms, buildings make decisions based on limited local observations, leading to suboptimal performance. For example, during summer peak hours, all buildings independently decide to pre-cool their spaces before peak pricing begins. This simultaneous cooling demand overloads the electrical grid, causing supply instability and higher costs for everyone.
@@
  • Lack of rigorous safety guarantees: Existing control methods lack rigorous safety guarantees [wang2025safe, ding2022safe] or rely on soft constraints, such as penalty terms in RL algorithms, to discourage unsafe behaviors. However, these methods cannot provide mathematical guarantees against constraint violations [brunke2022safe]. During the learning and exploration process, control systems may still take unsafe actions, such as exceeding battery capacity limits or causing equipment overloads. Moreover, when buildings operate as isolated islands with limited local observations and broken information connections, they are more likely to exhibit safety warnings and uncoordinated consumption patterns, leading to power grid-level equipment damage risks and overloads, resulting in significant economic losses.
@@
  • System complexity and algorithmic challenges: The inherent complexity of building energy systems poses significant algorithmic challenges for coordination [meng2024online, guo2022real]. As illustrated in Figure 1, the diverse building types and their complex interaction patterns create a challenging optimization landscape. Multi-objective optimization requirements involving comfort, efficiency, and safety lead to complex trade-offs that are difficult to balance simultaneously [Jang_Yan_Spangher_Spanos_2024]. The rapid development of distributed storage and renewable energy systems creates dynamic and uncertain environments with intermittent generation and variable demand patterns. These factors, combined with the instability of RL algorithms in high-dimensional continuous spaces, make convergence slow and solution quality unpredictable [MARLISA, MA-HVAC].
@@
To address these multi-agent RL challenges in building energy management, this paper proposes Spatial-Temporal Enhanced Safe Multi-Agent Coordination (STEMS), a novel spatial-temporal aware safe multi-agent reinforcement learning framework for coordinated building energy management. Our approach consists of two key components that work together to overcome the identified limitations. First, we develop a spatial-temporal graph representation learning framework that effectively captures the complex relationships among buildings. We propose a GCN-Transformer fusion architecture that processes both spatial dependencies through graph convolution and temporal patterns through attention mechanisms. This framework includes a selective information construction mechanism that enables intelligent collaboration among buildings by dynamically determining what information to share based on spatial-temporal contexts. The approach effectively handles high-dimensional state spaces through structured feature extraction and dimensionality reduction. Second, we design a safety-constrained multi-agent RL algorithm that provides mathematical guarantees for safe operation. We integrate Control Barrier Functions (CBFs) [brunke2022safe] as safety shields that ensure control actions remain within safe operating bounds. The algorithm incorporates a predictive safety checking mechanism that prevents unsafe actions from being executed while maintaining system performance. This enables precise safety control in continuous action spaces, addressing the critical need for reliable operation in real-world building energy systems. In summary, the main contributions of this work are threefold:
@@
  • We conduct extensive experiments on real-world datasets that validate the integration of spatial-temporal coordination and safety constraints enables STEMS to achieve 21% cost reduction over traditional methods, 65% safety violation reduction compared with learning-based methods, and maintains the lowest discomfort rate, demonstrating simultaneous optimization of energy efficiency, operational safety, and occupant comfort.
The remainder of this paper is organized as follows. Section II presents the system model and methodology, including problem formulation, the spatial-temporal graph representation learning framework, and the safety-constrained multi-agent RL algorithm. Section III presents experimental settings and comprehensive results on real-world building energy datasets. Section IV reviews related work, and Section V concludes the paper and discusses future research directions.
@@ II-A Problem Formulation
where $\mathcal{R}=\{R_{i}\}$ denotes the reward function set, $\mathcal{C}=\{C_{i}\}$ represents the constraint function set, and $\gamma$ is the discount factor.
@@ II-B Overall Framework
To simultaneously address spatial-temporal dependency modeling, safety constraint guarantees, and multi-objective optimization challenges in building energy management, STEMS adopts a modular design that integrates the spatial-temporal graph representation learning module, the CBF safety constraint checking module, and the multi-agent RL module. Figure 2 illustrates the overall architecture and information flow among these components. Initially, building agents collect comprehensive state observations $\mathcal{S}$ including internal device status, environmental conditions, and neighboring building information. These observations are then processed by the spatial-temporal graph representation learning module through observation embeddings, graph convolution and temporal attention layers to extract spatial-temporal features that capture inter-building relationships and temporal patterns. Subsequently multi-agent RL utilizes these enhanced spatial-temporal features to generate control actions $\mathcal{A}$ for each building through coordinated decision-making, where the rewards $\mathcal{R}$ and state transitions guide the continuous learning process.
@@
The renewable energy component maximizes renewable energy utilization:
@@
These four components work together to achieve balanced performance. The economic term minimizes electricity costs by optimizing consumption timing and renewable energy use. For system stability, the grid-level coordination with smooth building operations helps buildings work together to avoid overloads while maintaining steady power consumption. The comfort term allows flexible trade-offs between energy efficiency and occupant comfort based on real-time conditions and the renewable component maximizes green energy utilization across all buildings.
@@ II-D Safety-Constrained Decision Making
Building energy systems require strict safety constraints to prevent equipment damage and ensure reliable operation. Battery overcharging can cause thermal runaway and fire hazards. Power overloads can damage electrical equipment and cause grid instability. Our safety-constrained approach uses CBFs to provide mathematical safety guarantees under feasible conditions, unlike penalty-based methods that may still produce unsafe actions during training and exploration.
@@ II-D1 Control Barrier Function Framework
The battery state is updated by: $SOC_{t+1}^{i}=SOC_{t}^{i}+\frac{a_{battery}^{i}\cdot\Delta t}{\text{Capacity}_{i}}$.
Building Safety: Building power consumption cannot exceed capacity to prevent equipment overload:
@@
This ensures the executed action is as close as possible to the original policy output while maintaining safety. If it is feasible, the action is safe. Otherwise, emergency conservative actions including stopping battery charging and reducing HVAC power will be applied. This approach ensures reliable building energy management operations through mathematical safety guarantees.
@@ II-E Multi-Agent Training Algorithm
$\nabla_{\theta^{i}}J(\theta^{i})=\mathbb{E}_{s^{i},a^{i}\sim\mathcal{A}^{safe}_{i}}\left[\nabla_{\theta^{i}}\log\pi_{\theta^{i}}(a^{i}|s^{i})\cdot A^{i}(s^{i},a^{i})\right]$ (24)
where $A^{i}(s^{i},a^{i})=R^{i}+\gamma V_{\phi^{i}}(s^{\prime i})-V_{\phi^{i}}(s^{i})$ is the advantage function for agent $i$, and the expectation is taken only over the safe action space.
@@
$L(\phi^{i})=\mathbb{E}_{s^{i},a^{i}\sim\mathcal{A}^{safe}_{i}}\left[(R^{i}+\gamma V_{\phi^{i}}(s^{\prime i})-V_{\phi^{i}}(s^{i}))^{2}\right]$ (25)
@@
1:Initial policy parameters $\{\theta^{i}\}_{i=1}^{N}$, critic parameters $\{\phi^{i}\}_{i=1}^{N}$, spatial-temporal graph parameters $\psi$
2:Trained policy, critic and spatial-temporal graph networks
@@
13:  Update actor networks: $\theta^{i}\leftarrow\theta^{i}+\alpha_{\pi}\nabla_{\theta^{i}}J(\theta^{i})$
14:  Update critic networks: $\phi^{i}\leftarrow\phi^{i}+\alpha_{v}\nabla_{\phi^{i}}L(\phi^{i})$
15:  Update spatial-temporal graph parameters: $\psi\leftarrow\psi+\alpha_{g}\nabla_{\psi}L_{total}$
16:end for
17:return Trained policies $\{\pi_{\theta^{i}}^{*}\}_{i=1}^{N}$, critics $\{V_{\phi^{i}}\}_{i=1}^{N}$, and spatial-temporal graph networks $\psi$
@@ III-A Experimental Setup
TABLE I: Long-term Performance Comparison (12-Month Average Results).
Peak Consumption Rate Rate Viol. Rate
Rule-Based 1.000 1.000 1.000 1.000 1.000 0.130 0.351
MPC 0.872 0.914 0.983 0.975 0.981 0.654 0.330
Single-Agent SAC 0.824 0.867 0.856 0.813 0.925 0.485 0.223
MADDPG 0.795 0.817 0.854 0.856 0.936 0.357 0.197
MetaEMS 0.836 0.804 0.884 0.835 0.907 0.396 0.231
MARLISA 0.847 0.875 0.826 0.838 0.956 0.237 0.214
MADCQ 0.814 0.837 0.862 0.823 0.906 0.325 0.155
D-MAPPO 0.803 0.834 0.841 0.816 0.912 0.152 0.198
STEMS 0.792 0.824 0.821 0.805 0.883 0.132 0.056
[…]
@@ III-A1 Simulation Environment
We conduct experiments based on the CityLearn environment[citylearn], a multi-building energy management simulation platform that provides realistic building energy dynamics and grid interactions. The experimental dataset is derived from real building energy consumption data in Travis County, Texas. The dataset spans from August 2018 to August 2019, capturing a full year of operation under typical subtropical humid climate conditions: annual average temperature of 20.7°C, temperature range from -7.8°C to 42.8°C. This climate profile creates significant cooling-dominated energy demands during summer months and moderate heating requirements in winter, providing realistic conditions for evaluating multi-building energy coordination strategies. The data includes 1 hour interval measurements of electricity consumption, solar generation, battery storage, and environmental conditions. The experimental setup consists of 8 buildings in a neighborhood configuration with typical regional distribution (5 residential, 2 commercial, 1 mixed-use). The 8 buildings are randomly selected from the Travis County dataset based on above ratio to ensure representative diversity in building characteristics and energy consumption patterns. Most residential buildings show dual-peak patterns with morning (6-9h) and evening (18-22h) usage spikes, reflecting typical household activities with strong seasonal variations. Commercial buildings demonstrate business-hour concentrated loads (9-17h) with midday peaks around 12-13h. Mixed-use buildings combine both residential and commercial characteristics with variable all-day consumption patterns. Each building is equipped with controllable devices including battery energy storage systems (BESS), HVAC systems, solar photovoltaic panels, hot water tank. Buildings can communicate with each other to share state information and coordinate energy management decisions.
@@ III-A2 Baseline Methods
Single-Agent SAC: Independent Soft Actor-Critic reinforcement learning for each building without coordination[sac].
MADDPG: Multi-agent deep deterministic policy gradient with centralized training and decentralized execution[lowe2017multi-maddpg].
MetaEMS: Meta-learning based energy management system using model-agnostic meta-learning (MAML) framework[MetaEMS].
MARLISA: Multi-agent reinforcement learning with independent soft actor-critic operating independently[MARLISA].
MADCQ: Multi-agent deep constrained Q-learning with independent building-level optimization and heuristic constraint handling[saberi2024madcq].
D-MAPPO: Distributed multi-agent proximal policy optimization with soft penalty constraints for improved training stability[tariq2025dmappo].